OpenStudy (anonymous):

How do you set up the word problem "The measure of a supplement of an angle is 60 less than three times the measure of the complement of the angle. Find the measure of the angle."

OpenStudy (anonymous):

The supplement of an angle,\[(180{}^{\circ}-\theta ) \]is\[(180{}^{\circ}-\theta )\text{= }\]60 less\[(180{}^{\circ}-\theta )\text{= }-60{}^{\circ}\]than three times the complement of the angle.\[(180{}^{\circ}-\theta )\text{= }-60{}^{\circ}+3(90{}^{\circ}-\theta )\]\[\theta =15 {}^{\circ} \]